# An infinite family of pairs of distinct quartic Galois CM-fields with the same discriminant and regulator

### v1

@@ 1. Introduction
A number field is a subfield of $\mathbb{C}$ whose degree over $\mathbb{Q}$ is finite. Finding sufficient conditions for any two given number fields to be isomorphic or coincident is a fundamental problem. A famous result of such a problem is the Neukirch-Uchida theorem, which states that two number fields are isomorphic if their absolute Galois groups are isomorphic. Neukirch [11] proved the case when two number fields are Galois extensions over $\mathbb{Q}$, and then Uchida [16] proved the general case.
@@
We investigate the problem described in the beginning from the viewpoint of whether two number fields coincide if one or more of their invariants coincide. Specifically, we focus on the discriminant and the regulator because they are invariants of a number field appearing in the class number formula.
Throughout this paper, for a number field $K$, we denote by $\mathrm{disc}(K)$, $E(K)$, $\mathrm{reg}(K)$, $h(K)$ and $W(K)$ the discriminant, unit group, regulator, class number and the group of roots of unity contained in $K$, respectively.
@@
$\lim_{s\to 1}(s-1)\zeta_{K}(s)=\frac{2^{r_{1}}(2\pi)^{r_{2}}h(K)\mathrm{reg}(K)}{\lvert W(K)\rvert\sqrt{\lvert\mathrm{disc}(K)\rvert}},$
@@
    There is an infinite family of pairs of imaginary biquadratic fields $K_{1}$ and $K_{2}$ such that $\mathrm{disc}(K_{1})=\mathrm{disc}(K_{2})$ and $\mathrm{reg}(K_{1})=\mathrm{reg}(K_{2})=R$.
@@
    There is an infinite family of pairs of imaginary cyclic quartic fields $K_{1}$ and $K_{2}$ such that $\mathrm{disc}(K_{1})=\mathrm{disc}(K_{2})$ and $\mathrm{reg}(K_{1})=\mathrm{reg}(K_{2})=R$.
@@
The Hermite-Minkowski theorem states that for any integer $D$, there are only finitely many number fields with discriminant $D$.
@@ 2. Regulators of CM-fields
$Q\cdot\mathrm{reg}(K)=2^{r}\cdot\mathrm{reg}(K^{+}).$
@@
Theorem 2.1 (Nakamura [10], §2.4, Lemma 1).
If $\mathrm{disc}(K)/\mathrm{disc}(K^{+})^{2}$ does not divide $2^{n}$, then $Q=1$.
@@ 3. Regulators of imaginary biquadratic fields
For a square-free integer $m$, let $F_{m}=\mathbb{Q}(\sqrt{m})$. Let $m_{1}$ and $m_{2}$ be square-free positive integers such that $\gcd(m_{1},m_{2})=1$ and $m_{1}\equiv m_{2}\equiv 1\pmod{4}$.
Since $2m_{2}\equiv 2\pmod{4}$, we have $\mathrm{disc}(F_{2m_{2}})=8m_{2}$.
Since $-m_{1}\equiv 3\pmod{4}$, we have $\mathrm{disc}(F_{-m_{1}})=-4m_{1}$.
Since $-2m_{1}m_{2}\equiv 2\pmod{4}$, we have $\mathrm{disc}(F_{-2m_{1}m_{2}})=-8m_{1}m_{2}$.
Since $-2m_{1}\equiv 2\pmod{4}$, we have $\mathrm{disc}(F_{-2m_{1}})=-8m_{1}$.
Since $-m_{1}m_{2}\equiv 3\pmod{4}$, we have $\mathrm{disc}(F_{-m_{1}m_{2}})=-4m_{1}m_{2}$.
@@
$\mathrm{disc}(B_{-m_{1},2m_{2}})=\mathrm{disc}(F_{2m_{2}})\cdot\mathrm{disc}(F_{-m_{1}})\cdot\mathrm{disc}(F_{-2m_{1}m_{2}})=2^{8}m_{1}^{2}m_{2}^{2}.$
@@
$\mathrm{disc}(B_{-2m_{1},2m_{2}})=\mathrm{disc}(F_{2m_{2}})\cdot\mathrm{disc}(F_{-2m_{1}})\cdot\mathrm{disc}(F_{-m_{1}m_{2}})=2^{8}m_{1}^{2}m_{2}^{2}.$
@@
$\mathrm{disc}(B_{-m_{1},2m_{2}})=\mathrm{disc}(B_{-2m_{1},2m_{2}}).$
@@
$\frac{\mathrm{disc}(B_{-m_{1},2m_{2}})}{\mathrm{disc}(F_{2m_{2}})^{2}}=\frac{\mathrm{disc}(B_{-2m_{1},2m_{2}})}{\mathrm{disc}(F_{2m_{2}})^{2}}=2^{2}m_{1}^{2}.$
@@
$\mathrm{reg}(B_{-m_{1},2m_{2}})=\mathrm{reg}(B_{-2m_{1},2m_{2}})=2\cdot\mathrm{reg}(F_{2m_{2}}).$
@@ 4. Defining polynomials of cyclic quartic fields
±$\pm\sqrt{s(t^{2}+1)\pm s\sqrt{t^{2}+1}}.$
We denote by $K_{s,t}$ the splitting field of $f(s,t,X)$ over $\mathbb{Q}$.
When $K_{s,t}/\mathbb{Q}$ is a Galois extension, since all conjugate fields of $K_{s,t}$ over $\mathbb{Q}$ coincide, we obtain
$K_{s,t}=\mathbb{Q}\left(\!\sqrt{s(t^{2}+1)+s\sqrt{t^{2}+1}}\,\right).$
@@
To show that $K_{s,t}$ is a cyclic quartic field, we verify the conditions of Theorem 4.1 for $f(s,t,X)$.
Since $t^{2}+1\not\in\mathbb{Q}^{2}$, the polynomial $f(s,t,X)$ cannot be factored into a product of polynomials with rational coefficients of degree less than $4$. Thus, $f(s,t,X)$ is irreducible over $\mathbb{Q}$.
Again, since $t^{2}+1\not\in\mathbb{Q}^{2}$, we see that $\mathbb{Q}(\sqrt{t^{2}+1})$ is a quadratic field. Putting
@@
we obtain $K_{s,t}=\mathbb{Q}(\theta)$ and $\mathbb{Q}(\sqrt{t^{2}+1})=\mathbb{Q}(\theta^{2})$. Hence, $\mathbb{Q}(\sqrt{t^{2}+1})$ is a subfield of $K_{s,t}$. ∎
@@
Since both $u-t$ and $u+t$ are integers, we obtain
@@
In either case, we have $u-t=u+t$, which implies $t=0$. ∎
@@
By Lemma 4.2 and 4.3, $f(s,t,X)$ is irreducible over $\mathbb{Q}$, and hence $K_{s,t}$ is a cyclic quartic field. ∎
@@
Let $s$, $s^{\prime}$ and $t$ be non-zero rational numbers satisfying $ss^{\prime}\not\in\mathbb{Q}^{2}$. Then $K_{s,t}=K_{s^{\prime},t}$ if and only if $t^{2}+1\in ss^{\prime}\mathbb{Z}^{2}$.
@@
Hence, $\mathbb{Q}(\sqrt{ss^{\prime}})$ is a subfield of $K_{s,t}$. From the assumption $ss^{\prime}\not\in\mathbb{Q}^{2}$ of the lemma, $\mathbb{Q}(\sqrt{ss^{\prime}})$ is a quadratic field. By Lemma 4.2, $K_{s,t}$ contains $\mathbb{Q}(\sqrt{t^{2}+1})$. Since $K_{s,t}$ is cyclic, it has only one quadratic subfield. Therefore, $\mathbb{Q}(\sqrt{ss^{\prime}})=\mathbb{Q}(\sqrt{t^{2}+1})$ holds. Thus, we obtain $\mathbb{Q}(\sqrt{ss^{\prime}})=\mathbb{Q}(\sqrt{t^{2}+1})$, which implies that $t^{2}+1\in ss^{\prime}\mathbb{Q}^{2}$.
Conversely, suppose that $t^{2}+1\in ss^{\prime}\mathbb{Q}^{2}$. Then, $\mathbb{Q}(\sqrt{ss^{\prime}})=\mathbb{Q}(\sqrt{t^{2}+1})$. Since $ss^{\prime}\not\in\mathbb{Q}^{2}$ implies $t^{2}+1\not\in\mathbb{Q}^{2}$, by Lemma 4.2, we have $\sqrt{ss^{\prime}}\in K_{s,t}$. Therefore, $K_{s^{\prime},t}\subseteq K_{s,t}$. We can show the reverse inclusion in the same way. ∎
@@
Let $t$ be an integer with $t\equiv 3$ or $5\pmod{8}$. Then, $t^{2}+1\not\in 2\mathbb{Z}^{2}$.
@@
For the case $t\equiv 3\pmod{8}$, there is some $k\in\mathbb{Z}$ such that $t=8k+3$. If $t^{2}+1\in 2\mathbb{Z}^{2}$ then there is some $u\in\mathbb{Z}$ such that $t^{2}+1=2u^{2}$. Thus we obtain
$u^{2}=\frac{t^{2}+1}{2}=\frac{1+(8k+3)^{2}}{2}=32k^{2}+24k+5\equiv 5\pmod{8},$
which is impossible. Therefore, we must have $t^{2}+1\not\in 2\mathbb{Z}^{2}$.
We can show the case $t\equiv 5\pmod{8}$ in the same way. ∎
@@
Let $t$ be an integer with $t\equiv 3$ or $5\pmod{8}$. Then, $K_{s,t}\neq K_{2s,t}$.
@@
$ss^{\prime}\mathbb{Z}^{2}=2s^{2}\mathbb{Z}^{2}=2\mathbb{Z}^{2}.$
By Lemma 4.7, we have $t^{2}+1\not\in ss^{\prime}\mathbb{Z}^{2}$. By Lemma 4.5, we obtain the desired result. ∎
@@ 5. Discriminants of cyclic quartic fields
    For any odd prime $q$, if $q\nmid s$ and $q\nmid t^{2}+1$, then the $q$-exponent of the discriminant is 00.
@@
$x^{4}n^{2}=y^{2}\equiv 1,\quad m\equiv 2\pmod{8}.$

### v2

@@ 1. Introduction
A number field is a subfield of $\mathbb{C}$ whose degree over $\mathbb{Q}$ is finite. Finding sufficient conditions for any two given number fields to be isomorphic or coincident is a fundamental problem. A famous result of such a problem is the Neukirch-Uchida theorem, which states that two number fields are isomorphic if their absolute Galois groups are isomorphic as topological groups. Neukirch [13] proved the case when two number fields are Galois extensions over $\mathbb{Q}$, and then Uchida [18] proved the general case.
@@
We investigate the problem described at the beginning from the viewpoint of whether two number fields coincide if one or more of their invariants coincide. Specifically, we focus on the discriminant and the regulator because they are invariants of a number field appearing in the class number formula.
Throughout this paper, for a number field $K$, we denote by $\operatorname{disc}(K)$, $E(K)$, $\operatorname{reg}(K)$, $h(K)$ and $W(K)$ the discriminant, unit group, regulator, class number and the group of roots of unity contained in $K$, respectively.
@@
$\lim_{s\to 1}(s-1)\zeta_{K}(s)=\frac{2^{r_{1}}(2\pi)^{r_{2}}h(K)\operatorname{reg}(K)}{\lvert W(K)\rvert\sqrt{\lvert\operatorname{disc}(K)\rvert}}$,
@@
    There is an infinite family of pairs of distinct imaginary biquadratic fields $K_{1}$ and $K_{2}$ such that $\operatorname{disc}(K_{1})=\operatorname{disc}(K_{2})$ and $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})=R$.
@@
    There is an infinite family of pairs of distinct imaginary cyclic quartic fields $K_{1}$ and $K_{2}$ such that $\operatorname{disc}(K_{1})=\operatorname{disc}(K_{2})$ and $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})=R$.
@@
By the Hermite-Minkowski theorem, for any integer $D$, there are only finitely many number fields with discriminant $D$; hence the discriminant analogues of Theorems 1.2 (i) and 1.3 (i) cannot hold: there is no infinite family of pairwise distinct imaginary biquadratic fields or imaginary cyclic quartic fields having the same discriminant, and therefore in Theorems 1.2 (ii) and 1.3 (ii), the common discriminants of each pair are necessarily unbounded.
@@ 2. Regulators of CM-fields
$Q\cdot\operatorname{reg}(K)=2^{r}\cdot\operatorname{reg}(K^{+}).$
@@
Theorem 2.1 (Nakamula [12], §2.4, Lemma 1).
If $\operatorname{disc}(K)/\operatorname{disc}(K^{+})^{2}$ does not divide $2^{n}$, then $Q=1$.
@@ 3. Regulators of imaginary biquadratic fields
For a square-free integer $m$, let $F_{m}=\mathbb{Q}(\sqrt{m})$. Let $m_{1}$ and $m_{2}$ be square-free positive integers greater than $1$ such that $\gcd(m_{1},m_{2})=1$ and $m_{1}\equiv m_{2}\equiv 1\pmod{4}$.
Since $2m_{2}\equiv 2\pmod{4}$, we have $\operatorname{disc}(F_{2m_{2}})=8m_{2}$.
Since $m_{1}\equiv 3\pmod{4}$, we have $\operatorname{disc}(F_{-m_{1}})=-4m_{1}$.
Since $2m_{1}m_{2}\equiv 2\pmod{4}$, we have $\operatorname{disc}(F_{-2m_{1}m_{2}})=-8m_{1}m_{2}$.
Since $2m_{1}\equiv 2\pmod{4}$, we have $\operatorname{disc}(F_{-2m_{1}})=-8m_{1}$.
Since $m_{1}m_{2}\equiv 3\pmod{4}$, we have $\operatorname{disc}(F_{-m_{1}m_{2}})=-4m_{1}m_{2}$.
@@
$\operatorname{disc}(B_{-m_{1},2m_{2}})=\operatorname{disc}(F_{2m_{2}})\cdot\operatorname{disc}(F_{-m_{1}})\cdot\operatorname{disc}(F_{-2m_{1}m_{2}})=2^{8}m_{1}^{2}m_{2}^{2}.$
@@
$\operatorname{disc}(B_{-2m_{1},2m_{2}})=\operatorname{disc}(F_{2m_{2}})\cdot\operatorname{disc}(F_{-2m_{1}})\cdot\operatorname{disc}(F_{-m_{1}m_{2}})=2^{8}m_{1}^{2}m_{2}^{2}.$
@@
$\operatorname{disc}(B_{-m_{1},2m_{2}})=\operatorname{disc}(B_{-2m_{1},2m_{2}}).$
@@
$\frac{\operatorname{disc}(B_{-m_{1},2m_{2}})}{\operatorname{disc}(F_{2m_{2}})^{2}}=\frac{\operatorname{disc}(B_{-2m_{1},2m_{2}})}{\operatorname{disc}(F_{2m_{2}})^{2}}=2^{2}m_{1}^{2}.$
@@
$\operatorname{reg}(B_{-m_{1},2m_{2}})=\operatorname{reg}(B_{-2m_{1},2m_{2}})=2\cdot\operatorname{reg}(F_{2m_{2}}).$
@@ 4. Defining polynomials of cyclic quartic fields
$\pm\sqrt{s(t^{2}+1)\pm s\sqrt{t^{2}+1}}.$
We denote by $K_{s,t}$ the splitting field of $f(s,t,X)$ over $\mathbb{Q}$. Put
$\theta$ $=\sqrt{s(t^{2}+1)+s\sqrt{t^{2}+1}},$
$\theta^{\prime}$ $=\sqrt{s(t^{2}+1)-s\sqrt{t^{2}+1}}.$
Since
$\sqrt{t^{2}+1}=\frac{\theta^{2}}{s}-(t^{2}+1),$
$\theta^{2}\theta^{\prime 2}=s^{2}t^{2}(t^{2}+1),$
we have
$\theta^{\prime}=\pm\frac{st\sqrt{t^{2}+1}}{\theta}\in\mathbb{Q}(\theta).$
Therefore, all roots of $f(s,t,X)$ lie in $\mathbb{Q}(\theta)$, and we obtain
$K_{s,t}=\mathbb{Q}(\theta)=\mathbb{Q}\left(\!\sqrt{s(t^{2}+1)+s\sqrt{t^{2}+1}}\,\right).$
@@
To show that $K_{s,t}$ is a cyclic quartic field, we verify the conditions of Theorem 4.1 for $f(s,t,X)$. Put
$f(s,t,X)=X^{4}+aX^{2}+b$
with
$a=-2s(t^{2}+1),\quad b=s^{2}t^{2}(t^{2}+1).$
Since $b=(st)^{2}(t^{2}+1)$ and $t^{2}+1\not\in\mathbb{Q}^{2}$, we have $b\not\in(\mathbb{Q}^{\times})^{2}$. Moreover,
$a^{2}-4b=4s^{2}(t^{2}+1)^{2}-4s^{2}t^{2}(t^{2}+1)=4s^{2}(t^{2}+1),$
and hence $a^{2}-4b\not\in(\mathbb{Q}^{\times})^{2}$.
We prove that $f(s,t,X)$ is irreducible over $\mathbb{Q}$. Since $f(s,t,X)$ is an even polynomial, any factorization over $\mathbb{Q}$ into two quadratic factors must be of one of the following forms:
$(X^{2}+u)(X^{2}+v),\quad u,\,v\in\mathbb{Q},$
$(X^{2}+pX+q)(X^{2}-pX+q),\quad p,\,q\in\mathbb{Q}.$
In the first case, we have $a=u+v$ and $b=uv$, and hence $a^{2}-4b=(u-v)^{2}\in(\mathbb{Q}^{\times})^{2}$, which contradicts $a^{2}-4b\not\in(\mathbb{Q}^{\times})^{2}$. In the second case, the constant term is $q^{2}$, and hence $b=q^{2}\in(\mathbb{Q}^{\times})^{2}$, which contradicts $b\not\in(\mathbb{Q}^{\times})^{2}$. Therefore, $f(s,t,X)$ is irreducible over $\mathbb{Q}$.
Next, we compute
$b(a^{2}-4b)$ $=s^{2}t^{2}(t^{2}+1)\cdot 4s^{2}(t^{2}+1)$
$=4s^{4}t^{2}(t^{2}+1)^{2}$
$=\bigl(2s^{2}t(t^{2}+1)\bigr)^{2}\in(\mathbb{Q}^{\times})^{2}.$
Thus $f(s,t,X)$ satisfies the assumption of Theorem 4.1, and hence the field generated by a root of $f(s,t,X)$ is a cyclic quartic field. In particular, if we put
@@
then $K_{s,t}=\mathbb{Q}(\theta)$ (as shown above in this section), and therefore $K_{s,t}$ is a cyclic quartic field.
Finally, we have
$\sqrt{t^{2}+1}=\frac{\theta^{2}}{s}-(t^{2}+1)\in\mathbb{Q}(\theta)=K_{s,t},$
and hence $\mathbb{Q}(\sqrt{t^{2}+1})\subseteq K_{s,t}$. Since $t^{2}+1\not\in\mathbb{Q}^{2}$, the field $\mathbb{Q}(\sqrt{t^{2}+1})$ is a real quadratic field. ∎
@@
Since both utu-t and $u+t$ are integers, we obtain
@@
In either case, we have ut=u+tu-t=u+t, which implies $t=0$. ∎
@@
By Lemmas 4.2 and 4.3, $f(s,t,X)$ is irreducible over $\mathbb{Q}$, and hence $K_{s,t}$ is a cyclic quartic field. ∎
@@
Let $s$, $s^{\prime}$ and $t$ be non-zero rational numbers satisfying $ss^{\prime}\not\in\mathbb{Q}^{2}$ and $t^{2}+1\not\in\mathbb{Q}^{2}$. Then $K_{s,t}=K_{s^{\prime},t}$ if and only if $t^{2}+1\in ss^{\prime}\mathbb{Q}^{2}$.
@@
Hence, $\mathbb{Q}(\sqrt{ss^{\prime}})$ is a subfield of $K_{s,t}$. From the assumption $ss^{\prime}\not\in\mathbb{Q}^{2}$ of the lemma, $\mathbb{Q}(\sqrt{ss^{\prime}})$ is a quadratic field. By the assumption $t^{2}+1\not\in\mathbb{Q}^{2}$ and Lemma 4.2, $K_{s,t}$ contains $\mathbb{Q}(\sqrt{t^{2}+1})$. Since $K_{s,t}$ is cyclic, it has only one quadratic subfield. Therefore, we obtain $\mathbb{Q}(\sqrt{ss^{\prime}})=\mathbb{Q}(\sqrt{t^{2}+1})$, which implies that $t^{2}+1\in ss^{\prime}\mathbb{Q}^{2}$.
Conversely, suppose that $t^{2}+1\in ss^{\prime}\mathbb{Q}^{2}$. Since $t^{2}+1\not\in\mathbb{Q}^{2}$, by Lemma 4.2, we have $\sqrt{t^{2}+1}\in K_{s,t}$, and hence $\sqrt{ss^{\prime}}\in K_{s,t}$. Then we observe that
$\sqrt{\frac{s^{\prime}}{s}}=\frac{\sqrt{ss^{\prime}}}{s}\in K_{s,t}.$
It follows that
$\sqrt{s^{\prime}(t^{2}+1)+s^{\prime}\sqrt{t^{2}+1}}=\sqrt{\frac{s^{\prime}}{s}}\cdot\sqrt{s(t^{2}+1)+s\sqrt{t^{2}+1}}\in K_{s,t},$
which yields $K_{s^{\prime},t}\subseteq K_{s,t}$. We can show the reverse inclusion in the same way. ∎
@@
Let $t$ be an integer with $t\equiv 3$ or $5\pmod{8}$. Then, $t^{2}+1\not\in 2\mathbb{Q}^{2}$.
@@
Let $t$ be an integer with $t\equiv 3$ or $5\pmod{8}$. Suppose that $t^{2}+1\in 2\mathbb{Q}^{2}$. Then there exists $q\in\mathbb{Q}$ such that $t^{2}+1=2q^{2}$. Moreover, $q$ can be expressed as an irreducible fraction as follows:
$q=\frac{m}{n},\quad\gcd(m,n)=1,\quad n>0.$
Then we have
$t^{2}+1=2\cdot\left(\frac{m}{n}\right)^{2}.$
Multiplying both sides by $n^{2}$ and dividing by $2$, we obtain
$\frac{t^{2}+1}{2}\cdot n^{2}=m^{2}.$
Since $t$ is odd, $(t^{2}+1)/2$ on the left-hand side is an integer. If $n$ has a prime factor $p$, then $p^{2}\mid m^{2}$, implying $p\mid m$, which contradicts $\gcd(m,n)=1$. Hence $n=1$. Therefore,
$\frac{t^{2}+1}{2}=m^{2}.$
On the other hand, since $t\equiv 3$ or $5\pmod{8}$, we have $t^{2}\equiv 9\pmod{16}$. Thus,
$\frac{t^{2}+1}{2}\equiv\frac{10}{2}\equiv 5\pmod{8}$.
Hence,
$m^{2}\equiv 5\pmod{8}$.
However, the only possible values of a square of an integer modulo $8$ are 0, 11 and $4$, which is a contradiction. Therefore, we must have $t^{2}+1\not\in 2\mathbb{Q}^{2}$. ∎
@@
Let $s$ and $t$ be non-zero integers with $t\equiv 3$ or $5\pmod{8}$. Then, $K_{s,t}\neq K_{2s,t}$.
@@
$ss^{\prime}\mathbb{Q}^{2}=2s^{2}\mathbb{Q}^{2}=2\mathbb{Q}^{2}.$
By Lemma 4.7, we have $t^{2}+1\not\in ss^{\prime}\mathbb{Q}^{2}$. Since $s$ and $t$ are non-zero integers, by Lemmas 4.3 and 4.5, we obtain the desired result. ∎
@@ 5. Discriminants of cyclic quartic fields
    For any odd prime $q$, if $q\nmid s$ and $q\nmid t^{2}+1$, then the $q$-exponent of the discriminant is 0.
@@
$x^{4}n^{2}=y^{2}\equiv 1\pmod{8},\quad m\equiv 2\pmod{8}$.